# Bridging the Preference Gap between Retrievers and LLMs

### 13 Jan 2024 v1

@@ 4 $best=\varnothing$;
6       for $d\leftarrow d^{\text{retr.}}_{1}$ to $d^{\text{retr.}}_{n}$ where $d$ not in $d^{\text{silv.}}$ do
7             $seq$ = concat($d$, $d^{\text{silv.}}$);
8             $cur$ = $R(seq)$;
9             if (cur > silv) and (cur > prev) then
10                   $best$ = $d$;
11                   $prev$ = $cur$;
@@
13
14      if best != $\varnothing$ then
15             add $best$ to the end of $d^{\text{silv.}}$;
16             $silv=prev$;
17             $prev=0$;
18             $best=\varnothing$
@@ 20             Break;
22
Algorithm 1 Synthesis Silver Passage Sequence
Training. Training with the silver passage sequence is straightforward, achieved by applying cross-entropy loss.
@@ 4.2 Reinforcement Learning
Although supervised learning can already help training the bridge model, it is still ineffective (we can see using SL alone results in mixed performance compared to using a retriever alone in Table 5). This is attributed to sparse supervision and the lack of end-to-end training on downstream results.
To address these issues, we apply RL to further train the bridge model. RL does not limit the possible passage sequences (we only consider permutations or deletions in the silver passage sequence in SL. However, the optimal passage sequence might require more complex manipulations, such as repetition, which RL can accommodate) and give more supervision to find the best passages sequence. Using the performance of downstream task as the reward, the bridge model is trained in an end-to-end manner on the downstream task. Specifically, our task can be formulated as a RL problem as follows:
Reward is the performance of tasks against their ground-truth labels as we focus on labeled downstream tasks.
Policy Model is the bridge model that need to be trained.
Action Space is limitted to Passage IDs as we are interested in adapting the retrieved passages
Training. The reward objective can be optimized by any off-the-shelf RL algorithm, e.g., proximal policy optimization (PPO).
@@ 5.1 Datasets and Baselines
Datasets. We consider four datasets, ranging from popular QA datasets to less investigated personalized ones. We also include one dataset that contains private email conversations (Avocado Email), which is unlikely to be included in the LLM’s pre-training datasets. This will further help us investigate the effectiveness of our proposed BGM model, as the LLM will have to rely on the retrieved passages. The summary of statistics is given in Table 1.
Open-domain QA. We conduct evaluations on two open-domain QA: Natural Questions (NQ) Kwiatkowski et al. (2019) and HotpotQA Yang et al. (2018). They both consist of questions, answers collected from Wikipedia and the Web. HotpotQA is a multi-hop QA dataset which requires findings and reasoning over multiple passages to answer the question. The candidate passages are retrieved from WikiPedia pages
Personalized Generation. We follow Li et al. (2023) to construct the personalized generation datasets. Specifically, suppose a user is writing a document, which we call the current document. Given the immediate context $x$ and the user’s personal context, we aim the finish the document as if the user had completed it. The immediate context $x$ or the input to the personalized writing task, is defined as the title and the start of the current document. The candidate passages are retrieved from documents authored by this user in the past. This includes datasets from two domains: Avocado Email (Email) Oard et al. (2015) and Amazon Book (Book) Ni et al. (2019).
Figure 4: An example from the Book dataset. The query consists of the instruction, the document (i.e., review for Book) title, document product (there will be no product information for Email), document start. The target is to complete the text given the query.
Baselines. We consider the state-of-the-art baselines: (1) GTR Ni et al. (2021) a widely recognized retriever that operates independently of LLMs; (2) a variant of GTR, termed Random, in which the order of passages retrieved by GTR is randomized; (3) Point-wise score ranking (PSR) Izacard et al. (2022). This is a variant from Izacard et al. (2022) where we substitute the decoder with our LLM and apply point-wise scoring reranking. This approach can be regarded as utilizing a reranker as a bridge model, but lacking the capability for dynamic selection; and (4) Additionally, we include a non-retrieval baseline, Naive, where no retrieval is conducted to aid in the generation process.
#Training #Val. #Test Avg. # words
@@
Table 1: Statistic of the 4 datasets.
@@ 5.2 LLM and Hyperparamters
We select the T5-XXL (11B) Raffel et al. (2020b) model as our bridge model for most experiments. In the supervised learning stage, the T5-11B model is fine-tuned with a base learning rate of 0.001. A linear warmup scheduler is used for the first 1,000 training steps. Additionally, the square root normalized decay of the learning rate is applied. The model is trained until its performance converges on the validation set. Decoding is performed using beam search with a beam size of 4. We use the PaLM2 model Anil et al. (2023), a new state-of-the-art LLM, as our LLM. It adopts temperature sampling as the decoding strategy. The parameters of PaLM2 are frozen, and we set the temperature to 0 to make the output deterministic. We use Exact-Match (EM) and BLEU as the metric to select the best prompt among prompt variants. It is also used as the reward to train the bridge model in the RL stage. The “Top-K” in Eq. 2 is set to 5.
@@ 5.3 Evaluation Results and Analysis
Table 2: Performance of all 4 datasets
Superior of BGM. Table 2 reports the overall performance. We can see that
(1) The proposed BGM outperforms all the 4 baselines in all 4 datasets. This clearly indicates BGM is effective in adapting the retrieved passages.
(2) Compared to the Naive approach, BGM shows significant improvement overall. However, an exception is observed in the Book dataset, where BGM’s improvement is more modest. This suggests that the LLM already possesses a substantial amount of relevant knowledge, likely due to the inclusion of Amazon reviews (which the Book dataset comprises) in the LLM’s pre-training data. Consequently, retrieval is not always essential in this context. This also explains why the Naive approach outperforms other baselines (except BGM) on the Book dataset. It’s important to note that BGM still manages to show improvement, albeit to a lesser extent, in this scenario
(3) Compared to the GTR approach, BGM demonstrates substantial improvement. The NQ dataset shows the least improvement, as most instances only require one retrieved passage, which both GTR and BGM can successfully include. On the other hand, the HotpotQA dataset shows the most significant improvement, indicating that HotpotQA may be more sensitive to irrelevant passages. It’s noteworthy that Random and GTR perform similarly, while GTR shows improvement over Naive (with the exception of the Book dataset). This suggests that the ranking itself does not significantly impact performance, but the selection of passages does
(4) Compared to the PSR approach, BGM once again demonstrates significant improvement. This indicates that pure reranking alone is not sufficient for the bridge model. Selection must also be taken into account. It’s notable that PSR performs similarly to GTR, further suggesting that reranking alone does not greatly affect performance
@@
Table 3: Ablation results - threshold the PSR
@@
Table 4: Ablation results - different silver passage sequence for SL
@@
Table 5: Ablation results - SL only
@@
FLAN-T5-XXL (BGM) 45.37 35.64 10.42 12.07
Table 6: Ablation results - different bridge model size
@@
Data NQ HotpotQA Email Book
Metric EM EM BLEU BLEU
@@
Table 7: Ablation results - different LLM size
@@
Table 8: Ablation results - the generability of BGM across various datasets and LLM sizes.
Understanding BGM. We are interested in what happen when we optimize the BGM, whether BGM’s improvement is trival, and how some important setting affects the BGM.
@@
(1) Whether the selection can be achieved naively by manually threshold the PSR? We conducted an ablation experiment using various thresholds for PSR, as shown in Table 3. The term “Top-K” in rows 3 to 6 refers to the selection of only the Top-K passages from PSR reranked passages. The results are mixed, but consistently inferior to those of BGM. This suggests that a naive manual threshold applied to the reranking model is insufficient to meet our objectives. Therefore, it is necessary to consider both reranking and dynamic selection simultaneously
@@
(2) How different silver passage sequence will affect the final performance? In Table 2, we employed greedy search passage sequences as silver passage sequence for SL (Sec. 4.1). However, lacking ground truth passage sequences, we also explored whether this approach is superior to others. Our ablation experiment in Table 4 involved using various types of silver passage sequence. The first row, labeled ’GTR’, indicates the use of GTR’s retrieved passages as silver passage sequence. Similarly, ’PSR’ refers to using PSR reranked passages as silver passage sequence. In the final row, we used greedy search passage sequences, representing the current version of BGM. The results demonstrate that the quality of silver passage sequence significantly affects downstream task performance. While using PSR as silver passage sequence shows improvement over GTR on three datasets, Greedy (BGM) further enhances performance and achieves the best results. Identifying potentially better silver passage sequence is left for future work.
@@
(3) How helpful RL is to the final performance? BGM integrates SL and RL, and we aim to assess the effectiveness of each component. An ablation experiment is detailed in Table 5. We can observe that BGM, when operating with only SL, performs significantly worse than the full BGM model in NQ, HotpotQA, and Email, and in some cases, it even underperforms GTR. This suggests that SL alone is inadequate, highlighting the necessity of incorporating RL. It’s important to note that removing SL is not feasible, as this would result in an excessively large search space for RL to effectively learn.
@@
(4) How different size of bridge model affect the final performance? In Table 2, we utilized Flan-T5-XXL (11B) as the bridge model, which is already smaller than the LLM (24B). However, we are curious about the feasibility of using an even more lightweight language model (LM) as the bridge model. The ablation study shown in Table 6 presents the outcomes of employing bridge models of various sizes. It is evident that all three sizes (large, XL, and XXL) surpass the performance of the setup without a bridge model (i.e., GTR), with the largest size yielding the best results. This demonstrates that a bridge model is beneficial even at a smaller scale, and that larger sizes lead to further improvements.
@@
(5) How different size of LLMs affect the final performance? In Table 2, Palm2-S was used as the LLM. We are interested in evaluating the effectiveness of the bridge model with different sizes of LLMs. We conducted experiments using Palm2-XXS, as detailed in Table 7. It is important to note that the smaller LLM struggles with personalized generation datasets (i.e., Email and Book), resulting in BLEU scores lower than 1%. We opted not to report these results as they may not accurately reflect the trend. However, observations from NQ and HotpotQA suggest that BGM significantly outperforms all baselines by a large margin. This indicates that BGM is effective even with a smaller LLM.
@@
(6) Can a trained bridge model generalize to different datasets and LLMs? The effectiveness of the bridge model across various datasets and LLMs has been demonstrated. A more ambitious objective is to extend this performance to new contexts without additional training. We conducted experiments to investigate the generalizability of the bridge model. The upper section of Table 8 presents the data on generalizability across different datasets. Row 2 shows the performance of BGM (i.e., the bridge model) when trained exclusively on the NQ dataset and then tested on three other unseen datasets. Similarly, row 3 describes training BGM solely on the Email dataset and testing it on the other three. In all cases, the performance is inferior to that of BGM when both training and testing are conducted on the same datasets. This result is expected, as no specific techniques have been developed for dataset generalization, a topic we leave for future work.
@@
In the lower section of Table 8, we present the results of BGM when tested on Palm2-XXS, but trained on Palm2-S. Comparing with results of BGM both trained and tested on Palm2-XXS (as shown in row 1 of the bottom section), the mismatch between LLM sizes leads to a significant decline in performance. This suggests that BGM’s ability to generalize across different LLM sizes is currently limited. Addressing this limitation is considered an important direction for future research.
@@
(8) Case Studies. We provided examples of GTR, PSR, and BGM in Table 9 in the Appendix for the NQ dataset. For question I, both GTR and PSR yield the same incorrect output, even though their contexts have different orders. Only BGM provides the correct answer, indicating that additional irrelevant context can be noisy and detrimental to RAG’s performance. In question II, unlike question I, GTR and PSR produce different answers. This suggests that the ranking can sometimes influence the results. While the specific mechanism by which ranking affects outcomes remains unclear (we regard it as a valuable topic for future research), it is evident that we need to consider ranking, as it may or may not affect performance. Again, BGM selects the most relevant passage and delivers the correct answer. In question III, none of the candidate passages contain the answer (they discuss FaZe Clan and the number of subscribers, but do not identify who has the most subscribers). GTR and PSR provide incorrect answers, as their additional context is unhelpful. In contrast, BGM opts not to select any passages and answers the question using its own parameters, resulting in the correct answer. This demonstrates that retrieval-augmented processes are not always necessary, and BGM is capable of handling such cases

### 20 Feb 2024 v2

@@ 3 $R^{\text{best}}\leftarrow-\infty$;
5       for $c\leftarrow d^{\text{retr.}}_{1}$ to $d^{\text{retr.}}_{n}$ where $c\not\in d^{\text{silv.}}$ do
6             $d^{\prime}\leftarrow\mathtt{add\_to\_seq}(d^{\text{silv.}},c)$;
7             $R^{\text{cur}}\leftarrow R(d^{\prime})$;
8             if $R^{\text{cur}}>R^{\text{best}}$ then
9                   $c^{\text{best}}\leftarrow c$;
10                   $R^{\text{best}}\leftarrow R^{\text{cur}}$;
11
@@ 10                   $R^{\text{best}}\leftarrow R^{\text{cur}}$;
13      if $R^{\text{best}}>R^{\text{silv.}}$ then
14             $d^{\text{silv.}}\leftarrow\mathtt{add\_to\_seq}(d^{\text{silv.}},c^{\text{best}})$;
15             $R^{\text{silv}}\leftarrow R^{\text{best}}$;
16             $R^{\text{best}}\leftarrow-\infty$;
17
@@ 16             $R^{\text{best}}\leftarrow-\infty$;
Algorithm 1 Synthesis SPS
4.1 Supervised Learning
To conduct SL, the ground-truth passage sequence is required for each query. Existing approaches, which focus on obtaining the relevance score of each passage (see Section 1), are not applicable as we need to determine which combination of passages is most effective for the downstream tasks. To address this, we propose to synthesis silver passage sequence (SPS) by selecting only the useful passages. This is done by greedy search that incrementally selects the next passage that can maximize the downstream task performance.
Synthesising SPS using greedy search. We denote the downstream task performance when using a given passage sequence for RAG as $R(\cdot)$. For the edge case where the passage sequence is empty, $R(\varnothing)$ simply denotes the task performance without retrieval augmentation (i.e., no passage used). We start from an empty SPS $d^{\text{silv.}}=\varnothing$, and iteratively add the next best candidate passage to the sequence, measured based on the resulted task performance $R(d^{\text{silv.}})$. We stop until no improvement can be made to $R(d^{\text{silv.}})$. Algorithm 1 shows the pseudo-code for synthesising SPS. The training with the SPS is achieved by applying cross-entropy loss.
@@ 4.2 Reinforcement Learning
Although SL can already help training the bridge model, it is still ineffective – we observe using SL alone results in mixed performance (see Table 5). This is attributed to sparse supervision (see Sec. 1) and the lack of end-to-end training on downstream results.
To address these issues, we apply RL to continue the training of the bridge model. In SL, we only consider permutations or deletions in the SPS, while RL can accommodate more complex manipulations that an optimal passage sequence might require, such as repetition. Additionally, RL provides enhanced supervision beyond the silver sequences through the reward from sampled passage sequences. Using the performance of downstream task as the reward, the bridge model is trained in an end-to-end manner. Specifically, our task can be formulated as an RL problem – Reward is the performance of the downstream task, usually measured based on certain ground-truth labels, e.g., Extract Match or BLEU. Policy model is the bridge model that need to be trained. Action space is restricted to passage IDs (Sec. 3) as we are interested in organizing rather than revising the retrieved passages. In training, the reward objective can be optimized by any off-the-shelf RL algorithm, e.g., proximal policy optimization (PPO).
@@ 5.1 Datasets and Baselines
Datasets. We consider four datasets, ranging from popular QA datasets to personalized generation datasets. We also include one dataset that contains private email conversations (Avocado Email), which is unlikely to be included in the LLM’s pre-training datasets. This will further help us investigate the effectiveness of our proposed BGM model, as the LLM will have to rely on the retrieved passages. The summary of statistics is given in Table 1.
Open-domain QA. We conduct evaluations on two open-domain QA datasets: Natural Questions (NQ) Kwiatkowski et al. (2019) and HotpotQA Yang et al. (2018). They both consist of questions, answers collected from Wikipedia and the Web. HotpotQA is a multi-hop QA dataset which requires findings and reasoning over multiple passages to answer the question. The candidate passages are retrieved from Wikipedia pages.
Personalized Generation. We follow Li et al. (2023) to construct the personalized generation datasets. In this task, the query includes the start and necessary properties (e.g., title) of a document authored by a user. The objective is to complete the document as if it were finished by the same user. The candidate passages for this task are retrieved from documents previously authored by this user. This includes datasets from two domains: Avocado Email (Email) Oard et al. (2015) and Amazon Book (Book) Ni et al. (2019). An example from Book is given in Fig. 4.
Figure 4: An example from the Book dataset. The query consists of the instruction, title, product (there will be no product information for Email) and document start. The goal of the task is to generate the remaining part of the document.
Baselines. We consider the state-of-the-art baselines: (1) GTR Ni et al. (2021) a widely recognized retriever that operates independently of LLMs; (2) a variant of GTR, termed Random, in which the order of passages retrieved by GTR is randomized; (3) Point-wise score ranking (PSR) Izacard et al. (2022). This is a variant from Izacard et al. (2022) where we substitute the decoder with our LLM and apply perplexity distillation. This approach can be regarded as utilizing a reranker as a bridge model, but lacking the capability for dynamic selection; and (4) Additionally, we include a non-retrieval baseline, Naive, where no retrieval augmentation in the generation process.
#Training #Val. #Test Avg. #Tokens
@@
Table 1: Statistic of the 4 datasets. “Avg. # tokens” indicates the average number of tokens in the prompt (which includes the query and retrieved passages). We ensure that the length does not exceed the maximum length allowed by the LLM.
@@ 5.2 LLM and Hyperparamters
We select the T5-XXL (11B) Raffel et al. (2020b) model as our bridge model for most experiments. In the supervised learning stage, the T5-XXL model is fine-tuned with a base learning rate of 0.001. A linear warmup scheduler is used for the first 1,000 training steps. Additionally, the square root normalized decay of the learning rate is applied. The model is trained until its performance converges on the validation set. Decoding is performed using beam search with a beam size of 4. We use the PaLM2-S model Anil et al. (2023), a new state-of-the-art LLM, as our LLM. It adopts temperature sampling as the decoding strategy. The parameters of PaLM2-S are frozen, and we set the temperature to 0 to make the output deterministic. We use Exact-Match (EM) and BLEU metrics, depending on the dataset. It is also used as the reward to train the bridge model in the RL stage. The $K$ in “Top-K” in Eq. 2 is set to 5.
@@ 5.3 Evaluation Results and Analysis
Table 2: Performance of all 4 datasets.
Superiority of BGM. Table 2 reports the overall performance. We can see that
(1) The proposed BGM outperforms all the 4 baselines in all 4 datasets. This clearly indicates BGM is effective in adapting the retrieved passages. It’s noteworthy that Random and GTR perform similarly, suggesting that ranking has less impact than selection of passages does.
(2) Compared to the Naive approach, BGM shows significant improvement overall. An exception is observed in the Book dataset, where the improvement of BGM is less pronounced. This suggests that retrieval is not always essential in this context. This also explains why the Naive approach outperforms other baselines (except BGM) on the Book dataset. It’s important to note that BGM still manages to show improvement by dynamically deciding whether and how many passages to select. For example, in the Book dataset, for around half of the samples (approximately 25,000), BGM selects no passages for augmentation.
(3) Compared to the GTR approach, BGM demonstrates substantial improvement. Compared to HotpotQA, NQ shows smaller improvement, as most instances need only one passage, which both GTR and BGM can successfully include. Conversely, the HotpotQA shows a more substantial improvement, indicating that HotpotQA may be more sensitive to irrelevant passages.
(4) Compared to the PSR approach, BGM again demonstrates significant improvement. This indicates that pure reranking alone is not sufficient for the bridge model. Selection must also be taken into account. It’s notable that PSR performs similarly to GTR, further suggesting that reranking alone has limited impact.
@@
Table 3: Ablation - threshold the PSR.
@@
Table 4: Ablation - different SPS for SL.
@@
Table 5: Ablation - SL only.
@@
FLAN-T5-XXL 45.37 35.64 10.42 12.07
Table 6: Ablation - different BGM bridge model size.
@@
Data NQ HotpotQA NQ HotpotQA
Metric EM EM EM EM
@@
Table 7: Ablation - different LLM size.
@@
Table 8: Ablation - the generability of BGM across various datasets and LLM sizes.
Understanding BGM. Next, we study a few research questions about BGM and RAG.
@@
(1) Can we perform effectively passage selection for RAG by simply thresholding PSR? We conducted an ablation experiment using various thresholds for PSR, as shown in Table 3. “Top-K” in rows 3 to 6 denotes selecting only the top-k passages from PSR reranked passages. The results vary but are consistently lower than BGM’s. This suggests that a naive manual threshold applied to the reranking model is insufficient to achieve the desired objectives. Therefore, it is necessary to consider both reranking and dynamic selection simultaneously.
@@
(2) How does different SPS (sliver passage sequence) used for SL affect BGM performance? In Table 2, we employed greedy search passage sequences as SPS for SL (Sec. 4.1). Here, we explore whether this approach is superior to others. Our ablation experiment in Table 4 involved using various types of SPS. The first row, labeled “GTR”, indicates the use of GTR’s retrieved passages as SPS. Similarly, “PSR” refers to using PSR reranked passages as SPS. In the final row, we used greedy search passage sequences, representing the proposed version of BGM. The results demonstrate that the quality of SPS significantly affects downstream task performance. While using PSR as SPS shows improvement over GTR on three datasets, Greedy (BGM) further enhances performance and achieves the best results. Identifying potentially better SPS is left for future work.
@@
(3) How helpful is RL to BGM? BGM integrates SL and RL, and we aim to assess the effectiveness of each component. An ablation experiment is detailed in Table 5. We can observe that BGM, when operating with only SL, performs significantly worse than the full BGM model in NQ, HotpotQA, and Email, and in some cases, it even underperforms GTR. This suggests that SL alone is inadequate, highlighting the necessity of incorporating RL. Note that removing SL is likely ineffective, leading to a poorly initialized policy model for RL with an excessively large search space.
@@
(4) How does the size of bridge model affect the final performance? In Table 2, we utilized Flan-T5-XXL (11B) as the bridge model, which is already smaller than the PaLM2-S. However, we are curious about the feasibility of using an even more lightweight LM as the bridge model. The ablation study shown in Table 6 presents the outcomes of employing bridge models of various sizes. It is evident that all three sizes (large, XL, and XXL) surpass the performance of the setup without a bridge model (i.e., GTR), with the largest size yielding the best results. This demonstrates that a bridge model is beneficial even at a smaller scale, and that larger sizes lead to further improvements.
@@
(5) How does the size of LLMs affect the final performance? In Table 2, Palm2-S was used as the LLM. We are interested in evaluating the effectiveness of the bridge model with different sizes of LLMs. We conducted experiments using Palm2-XXS in Table 7. It is important to note that the smaller LLM struggles with personalized generation datasets (i.e., Email and Book), resulting in BLEU scores lower than 1%. We opted not to report these results as they may not accurately reflect the trend. However, observations from NQ and HotpotQA suggest that BGM significantly outperforms all baselines by a large margin. This indicates that BGM is effective even with a smaller LLM.
@@
(6) Can a bridge model generalize to different datasets and LLMs? Our experiments above demonstrate the effectiveness of the bridge model when trained “in-domain” (i.e., trained and tested on the same dataset) using PaLM2-S as the LLM. A more ambitious goal is to extend this performance to new datasets or LLMs without extra training. We conducted ablation experiments to investigate bridge model generalizability: Table 8’s upper section shows the results across different datasets. Row 2 shows the performance of the bridge model when trained exclusively on the NQ dataset and then tested on three other unseen datasets. Similarly, row 3 shows training BGM solely on the Email dataset and testing it on the other three. In all cases, performance falls short of BGM’s when trained and tested on the same datasets. This is expected, given the lack of techniques for dataset generalization, a topic we leave for future work.
@@
In the lower section of Table 8, we present the results of BGM when tested on Palm2-XXS, but trained on Palm2-S. Comparing with results of BGM both trained and tested on Palm2-XXS (see row 1 of the bottom section), the mismatch between the training and testing LLMs leads to a significant decline in performance. This suggests that BGM’s ability to generalize across different LLMs is currently limited. Addressing this is considered an important direction for future research.
@@
(7) Case Studies. We provided examples of GTR, PSR, and BGM in Table 9 in Appendix for the NQ dataset. For question I, both GTR and PSR yield the same incorrect output, even though both include a relevant passage for RAG. Only BGM provides the correct answer, indicating that additional irrelevant context can be noisy and detrimental to RAG’s performance. In question II, none of the candidate passages contain the answer (they discuss FaZe Clan and the number of subscribers, but do not identify who has the most subscribers). GTR and PSR provide incorrect answers, as their additional context is unhelpful. In contrast, BGM opts not to select any passages and answers the question using its own memory, resulting in the correct answer. This demonstrates that retrieval-augmented processes are not always necessary, and BGM is capable of handling such cases.